OpenStudy (anonymous):

5^(3x-2)=5^(2x+1) Can someone PLZ explain how to solve this?

hero (hero):

Since the bases are the same, 3x-2 = 2x+1

hero (hero):

Solve for x

OpenStudy (waleed_imtiaz):

if bases are the same, then their exponents must be equal... DO this and solve for x..... :) 3x-2=2x+1......x=3

hero (hero):

@waleed_imtiaz at least give her a chance to solve it on her own :/

OpenStudy (anonymous):

AHHH that makes sense now!! i got x=3

OpenStudy (anonymous):

THX!!

OpenStudy (waleed_imtiaz):

@Hero she solved the question...... My pleasure @joyce153